OpenStudy (anonymous):

Care to start me off on this integral? Or give me some tips/guidance? (2x - 1)/(x^2 + 2x +5)

OpenStudy (anonymous):

First step would be to separate it into two separate fractions (one that you can use the whole denominator as a u-sub)

OpenStudy (anonymous):

\[ zero \ for \ 2x-1 \ = x = - \frac {1} {2} \\ f( - \frac {1} {2}) = - \frac {1} {2} ^2 + 2 (- \frac {1} {2} ) + 5 \\ \ \ \ = \frac {1}{4} - 1 + 5 \\ = \frac { 1 - ( 1*4) + ( 4*5) } {4} \\ = \ \frac {23}{4}\]

OpenStudy (anonymous):

So.. (2x+2)/(x^2 + 2x +5) - 3/(x^2 + 2x +5)

OpenStudy (anonymous):

>-1/2

OpenStudy (anonymous):

Ooooookay..... two/three people didn't read the problem. I thiiiiiiink I can get it from what you've said, redham..... thanks.